A ball is thrown straight up at 17 m/s. How fast is the ball moving when it is 6.2 m above its release point?
Answer #1

Here ,

let the speed is v

Using third equation of motion

v^2 - u^2 = 2 * g h

v^2 - 17^2 = -2 * 9.8 * 6.2

solving for v
v = 12.94 m/s

the speed of the ball is 12.94 m/s